Make sure everything is explained please, i'll rate for sure Let f(1) = 123, 051<1 1, 1<r <2 a) Consider the even periodic extension, of period T = 4, of f(x). Sketch 3 periods, on the i1, 0<x< 1, 2, 1<x< 2. a) Consider the even periodic extension, of period T = 4, of f(x). Sketch 3 periods, on the interval -6
Answer #1

F(20) 36 3 ocaeci 1 14202 flae) 2 Griven T= 4 y 5 period. >> → fl4+20) = f(ze) flT+2) flat) → (4+0) 3 > o = 428 zeci f(ae) 1